OpenStudy (dj3strella):

Help me please 5. Given the equation AB + CD = 25, what conclusion can be drawn based on the Subtraction Property of Equality? AB CD = 25 AB = 25 CD AB = CD + 25 AB = CD 25 I think its D

OpenStudy (dj3strella):

@bibby

OpenStudy (studygurl14):

B

OpenStudy (bibby):

assuming b is AB = 25 - CD what studygurl said

OpenStudy (texaschic101):

AB + CD = 25 Ab = 25 - CD or CD = 25 - AB

OpenStudy (texaschic101):

typo...I didn't mean to put a little b...lol

OpenStudy (dj3strella):

Thank you everyone